# Towards Foundation Model for Chemical Reactor Modeling: Meta-Learning with Physics-Informed Adaptation

### v2

@@ Abstract
In this work, we present a novel application of foundation models for chemical reactor modeling. Accurate modeling of real-world chemical reactors through first-principles is often challenging, and the process of rebuilding and retraining models for each new chemical process is inefficient. This raises a critical question: can we develop a single, universal neural network (i.e., a foundation model) that can rapidly adapt to any new chemical process in a reactor? To address this, we propose a foundation model for chemical reactor modeling that employs a meta-learning approach, followed by physics-informed fine-tuning on new tasks with only a few data samples. Our model is designed to generalize across three classic reactor types: continuous stirred tank reactors, batch reactors, and plug flow reactors. Compared to conventional methods such as data-driven learning, physics-informed learning, transfer learning, and meta-learning, our approach demonstrates superior performance in few-shot scenarios. Specifically, it shows rapid adaptation to unseen reactions with varying integer orders across different reactor set-ups, requiring minimal data for fine-tuning.
@@ 1 Introduction
Research on foundation models has made significant progress in recent years. The term “foundation model” refers to any model trained on broad data, capable of adaptation (e.g., fine-tuning) to a broad spectrum of downstream tasks (Bommasani et al., 2021). The success of foundation models in computer science, such as ChatGPT in natural language processing (Achiam et al., 2023) and Sora in computer vision (Brooks et al., 2024), has spurred their applications in traditional engineering disciplines such as material science (Takeda et al., 2023; Miret and Krishnan, 2024) and biomolecular engineering (i.e., predict molecular properties using Simplified Molecular Input Line Entry System (SMILES) representation) (Ahmad et al., 2022; Yang et al., 2024), natural sciences such as astronomy (Lanusse et al., 2023), physics (McCabe et al., 2023; Herde et al., 2024) and biology (Bryant et al., 2022; Hayes et al., 2024), and scientific machine learning (Hao et al., 2024).
@@
However, research on foundation models in chemical engineering remains limited but is increasingly recognized as critical (Decardi-Nelson et al., 2024). While some progress has been made in developing foundation models for chemistry-related scientific questions (Horawalavithana et al., 2022) and for solving chemical and biological tasks such as biomedical question answering and molecular synthesis (Livne et al., 2023), much of this work frames the problem through a linguistic lens, primarily constructing language models for domain-specific question answering. To the best of our knowledge, no research has yet addressed the development of foundation models for chemical reactor modeling. To address this gap, we aim to develop a foundation model specifically designed for modeling a wide range of chemical processes in classical reactors. Our objective is to tackle the challenges of effectively and efficiently modeling nonlinear chemical processes in these reactors.
@@
Dynamic process modeling, particularly for chemical reactions in real-world reactors, plays a crucial role in applications such as process monitoring, optimization, and advanced process control (Wu et al., 2019a; Hassanpour et al., 2020; Ellis and Chinde, 2020; Jalanko et al., 2021; Li et al., 2024a). Three idealized models are used to estimate the most important process variables of different chemical reactors: continuous stirred tank reactors (CSTRs), batch reactors (BRs), and plug flow reactors (PFRs). All three types of chemical reactors can be described using a set of ordinary differential equations (ODEs) or partial differential equations (PDEs), commonly referred to as first-principles models (Fogler, 2010). However, the parameters of these first-principles models vary depending on the specific reactions and reactors involved, and they can be difficult to determine accurately in real-world applications. Thus, there has been a transition to utilize data-driven methods, such as machine learning-based models, such as machine learning-based models, to capture the system dynamics of chemical reactions. These methods include least squares regression (Bhadriraju et al., 2019; Tang, 2023), feedforward neural networks (Liu et al., 2024), recurrent neural networks (RNNs) (Wu et al., 2019a, b), universal differential equation-based hybrid models (Bangi et al., 2022), physics-informed RNNs (PIRNNs) (Zheng et al., 2023), neural ordinary differential equations (Sorourifar et al., 2023), and input convex neural networks (Wang et al., 2025). However, a significant hurdle in employing deep learning within scientific domains is the scarcity of experimental data, which is normally costly and time-consuming to acquire via physical experimentation (e.g., it is resource-intensive to collect data and build neural network models for every new chemical reaction in the reactor). To mitigate this challenge, transfer learning-based approaches have been suggested (Xiao et al., 2023, 2024). However, for transfer learning-based methods, the process of identifying analogous reactions with adequate data, pre-training, and then transferring to the designated reaction must be repeated for each new reaction with a considerable number of samples, typically hundreds or thousands (Zhuang et al., 2020).
This prompts the question: what if we develop one universal neural network capable of swiftly adapting to any new reaction? Specifically, our goal is to train a foundation model on diverse datasets comprising various chemical reactions in CSTRs, BRs, and PFRs. The model will then be fine-tuned with a minimal number of samples from new tasks, enabling rapid adaptation to specific chemical reactions for reactor modeling. The high-level system architecture is shown in Fig. 1.
Figure 1: High-level system architecture of foundation model for various chemical reactors.
@@
To achieve this goal, we employ a meta-learning technique to build the foundation model, specifically Reptile (Nichol and Schulman, 2018). After acquiring the foundation model through Reptile, a conventional method to adapt the model to any new processes is via few-shot adaptation (the term “shot” refers to the number of data samples used to fine-tune the neural network). However, conventional data-driven adaptation in few-shot settings may result in poor generalization for nonlinear systems, due to the complexity of their state-space dynamics. To enhance the generalizability of the foundation model and reduce the data required for adaptation, we draw inspiration from the success of physics-informed neural networks (Raissi et al., 2019). Specifically, we adopt a physics-informed adaptation approach by integrating domain knowledge, such as governing equations (e.g., mass and energy balances), into the adaptation process. Simulation data of hundreds of different reactions in three classical reactors is used to develop the foundation model that will be further adapted to new reaction processes with limited data available. It is demonstrated that the adaptation process from the foundation model to a specific reaction is efficient, requiring only a few shots for successful adaptation. In summary, this work develops a foundation model for chemical reactor modeling, which enables few-shot adaptation to any unseen chemical processes in the generic chemical reactors.
@@
The rest of this article is structured as follows: Section 2 reviews the current works on physics-informed learning and meta-learning. Section 3 presents the algorithms for the Reptile-based foundation model with physics-informed adaptation. Section 4 outlines the formulation of CSTRs, BRs, and PFRs, covering reactor modeling and the design of physics-informed loss functions. Section 5 assesses the few-shot performance of the proposed method on unseen fixed-order CSTRs, BRs, and PFRs. Section 6 extends the method to incorporate random integer reaction orders. Finally, Section 7 addresses the limitations of the current design and proposes directions for future research.
@@
The effectiveness of transfer learning heavily relies on the similarity between the data distributions of the target and source tasks. In practice, finding a source reaction closely aligned with the target reaction is often challenging. A common approach to address this challenge involves conducting a few experiments under various conditions to gather data and develop a first-principles model through parameter optimization via regression. Once this model is established, it can be used to simulate a larger dataset, which then serves to pretrain a neural network for transfer learning, covering the entire operational range of the target reaction. However, transfer learning that depends on creating first-principles models to generate simulation datasets for each new reaction can be time-consuming, often taking hours or even days, and its effectiveness is highly contingent on the accuracy of the developed first-principles model. In contrast, our foundation model delivers high performance within seconds, eliminating the need for dependencies on first-principles models.
@@ 2 Meta-Learning and Physics-Informed Learning
Our goal and approach differ from these works. Due to the nature of our tasks, we use an RNN as the base model to develop a foundational model for generic chemical reactors (i.e., different parameterized ODE systems). Our goal is not to solve the ODEs, as PINNs do (Nascimento et al., 2020), but to predict future states of the system based on the current state. Unlike conventional PINNs, which involve multiple applications of automatic differentiation, we fine-tune the meta-learned RNN using a physics-informed loss term computed with numerical methods, such as the forward Euler method, which requires fewer computational resources. Furthermore, multi-head PINNs (Zou and Karniadakis, 2023) may be too complex (e.g., too many heads) for our tasks, as we are considering theoretically countless reactions in different reactors due to various parameter combinations. This complexity makes it challenging to incorporate multi-head PINNs into existing industrial engineering frameworks for downstream real-time applications, such as model predictive control (MPC), whereas a simpler RNN is more suitable for our purposes. Furthermore, unlike existing foundation models such as ChatGPT (Achiam et al., 2023), our approach tackles the problem through meta-learning and physics-informed learning instead of reinforcement learning from human feedback or self-supervised learning. Furthermore, we believe that a successful foundation model should support both few-shot learning and continual learning, and existing literature indicates that meta-learning has significant potential for enabling continual learning (Riemer et al., 2018; Javed and White, 2019; Wang et al., 2024).
@@ 3.1 Class of Systems
In this work, we consider chemical processes that can be described by the following class of nonlinear ODEs:
@@ 3.2 Reptile for Meta-Learning
We begin by introducing the meta-learning algorithm used to build the foundation model. Specifically, we employ Reptile (Nichol and Schulman, 2018), a simplified version of MAML that prioritizes computational efficiency (Nichol et al., 2018). Unlike MAML, which relies on second-order derivatives, Reptile uses only first-order derivatives for meta-learning updates, making it more suitable for real-world applications. The primary goal of meta-training is to expose the model to a wide range of tasks, enabling it to learn initial parameters that facilitate rapid adaptation to new, unseen tasks. During this phase, Reptile optimizes the model’s parameters for efficient adaptation to different tasks with minimal updates. In essence, Reptile guides the model toward a solution $\theta$ that closely approximates the manifold of optimal solutions for each task $\tau$, by minimizing the Euclidean distance through repeated task sampling, model updates, and parameter adjustments. Mathematically, Reptile identifies network parameters $\theta$ that minimize the Euclidean distance $\mathcal{D}(\theta,\mathcal{W}_{\tau})$ for all tasks (i.e., locating a point near all solution manifolds), where $\mathcal{W}_{\tau}$ is the set of optimal parameters for task $\tau$:
@@ 3.3 Reptile with Physics-Informed Adaptation
Next, we develop a foundation model to learn the dynamics of various chemical reactions in three types of reactors. Our methodology comprises two distinct phases: meta-training and meta-testing. During meta-training (i.e., Algorithm 1), we immerse the model (i.e., specifically, an RNN) in a diverse array of chemical reactions across various reactor types. Specifically, the meta-training process involves two steps. First, we obtain the optimal parameters $\mathcal{W}_{\tau}$ for a specific task $\tau$ using the mean squared error (MSE) and the Adam optimizer (Kingma, 2014). Once $\mathcal{W}_{\tau}$ is determined, we update the network parameters $\theta$ using the Reptile update rule:
$\theta\leftarrow\theta+\epsilon(\mathcal{W}_{\tau}-\theta)$ (3)
where $\epsilon$ is the meta-optimization step size hyperparameter. This process is then repeated recursively for all tasks.
@@
5:        Let $\widetilde{\theta}\leftarrow\theta$.
@@
9:           Update $\widetilde{\theta}\leftarrow Adam(\mathcal{L}_{\tau_{i}}(\widetilde{y_{i}},y_{i}),\widetilde{\theta})$.
@@
12:        Update $\theta\leftarrow\theta+\alpha(\widetilde{\theta}-\theta)$.
@@
Algorithm 1 Reptile-Based Foundation Model Meta-Training
In the subsequent meta-testing phase (i.e., Algorithm 2), we select previously unseen chemical reactions and undertake physics-informed few-shot adaptations to them, respectively, leveraging the Reptile-based foundation model obtained during meta-training. It should be pointed out that we use few-shot data $x$ (i.e., $x\in\mathcal{X}\subset D$, where $\mathcal{X}$ denotes the set of few-shot training data) to compute the data-driven loss term $\mathcal{L}_{d}$ and use collocation points $x_{c}$ (i.e., $x_{c}\in\mathcal{X}_{c}\subset D$, where $\mathcal{X}_{c}$ denotes the set of collocation data points, and $\mathcal{X}\cap\mathcal{X}_{c}=\emptyset$) to compute the physics-informed loss term $\mathcal{L}_{p}$ (a collocation point refers to a specific location within the domain of interest where the governing physics equations are enforced as part of the training process). As we do not utilize any labeled output data for $\mathcal{L}_{p}$, there is no need to perform physical experiments to gather additional data, which remains consistent with our few-shot setting. Moreover, it should be noted that we use estimated parameter values in $\mathcal{L}_{p}$ (i.e., fixed values) instead of their ground truth during adaptation (this differs from PIRNNs developed in our previous work (Zheng et al., 2023), which use ground truth parameter values), as the explicit parameter values of the first-principles models are not available in real-world scenarios (if we have the first-principles model with true parameter values, there is no need to perform neural network modeling). The detailed formulation of physics-informed loss functions with respect to CSTRs, BRs, and PFRs will be presented in the following section, along with the parameterized first-principles models of CSTRs, BRs, and PFRs.
@@
Require: $M$: pre-trained foundation model
@@
We observed that employing the aforementioned approach to build a foundation model specifically for one type of reactor (e.g., CSTRs) can achieve a testing MSE on the order of $10^{-4}$ for normalized data easily with around 5 shots using a 2-layer RNN. However, based on our previous works (Wu et al., 2019a, b; Wang et al., 2025), we observed that a testing MSE on the order of $\mathbf{10^{-3}}$ for normalized data is indicative of satisfactory performance, rendering the model suitable for applications in process systems engineering such as neural network-based MPC.
@@
While developing a foundation model for each type of reactor might yield better performance for that specific reactor, a unified foundation model for all types of reactors could offer a significant advantage in generalizability. Specifically, it can adapt to the same chemical reaction across different reactor types, which a single-type foundation model cannot achieve. This adaptability is crucial when the real-world system involves the same reaction but in a different reactor, as a single-type model lacks information about other reactor types. This motivates us to develop a unified foundation model that can handle all three types of chemical reactors, rather than creating separate models for each type (our goal is not to develop an extremely precise model).
@@
4 Modeling of Chemical Reactors and Development of Physics-Informed Loss Functions

### v3

@@ Abstract
Developing accurate models for chemical reactors is often challenging due to the complexity of reaction kinetics and process dynamics. Traditional approaches require retraining models for each new system, limiting generalizability and efficiency. In this work, we take a step toward foundation models for chemical reactor modeling by introducing a neural network framework that generalizes across diverse reactor types and rapidly adapts to new chemical processes. Our approach leverages meta-learning to pretrain the model on a broad set of reactor dynamics, enabling efficient adaptation to unseen reactions with minimal data. To further enhance generalizability, we incorporate physics-informed fine-tuning, ensuring physically consistent adaptation to new reactor conditions. Our framework is evaluated across three integer-order fundamental reactor types - continuous stirred tank reactors, batch reactors, and plug flow reactors - demonstrating superior few-shot adaptation compared to conventional data-driven, physics-informed, and transfer learning approaches. By combining meta-learning with physics-informed adaptation, this work lays the foundation for a generalizable modeling framework, advancing the development of foundation models for chemical engineering applications.
@@ 1 Introduction
Research on foundation models has made significant progress in recent years. The term “foundation model” refers to any data-driven model trained on broad data, capable of adaptation (e.g., fine-tuning) to a broad spectrum of downstream tasks (Bommasani et al., 2021). The success of foundation models in computer science, such as ChatGPT in natural language processing (Achiam et al., 2023) and Sora in computer vision (Brooks et al., 2024), has spurred their applications in traditional engineering disciplines such as material science (Takeda et al., 2023; Miret and Krishnan, 2024) and biomolecular engineering (i.e., predict molecular properties using Simplified Molecular Input Line Entry System (SMILES) representation) (Ahmad et al., 2022; Yang et al., 2024), natural sciences such as astronomy (Lanusse et al., 2023), physics (McCabe et al., 2023; Herde et al., 2024) and biology (Bryant et al., 2022; Hayes et al., 2024), and scientific machine learning (Hao et al., 2024).
@@
However, research on foundation models in chemical engineering remains limited but is increasingly recognized as critical (Decardi-Nelson et al., 2024). While some progress has been made in developing foundation models for chemistry-related scientific questions (Horawalavithana et al., 2022) and for solving chemical and biological tasks such as biomedical question answering and molecular synthesis (Livne et al., 2023), much of this work frames the problem through a linguistic lens, primarily constructing language models for domain-specific question answering. To the best of our knowledge, no research has yet addressed the development of foundation models for chemical reactor modeling. To bridge this gap, we take a step toward developing a foundation model for chemical reactor modeling, designed to generalize across a wide range of chemical processes in classical reactor systems. Our goal is to enable efficient and accurate modeling of nonlinear reaction dynamics while reducing the need for system-specific retraining.
@@
Dynamic process modeling, particularly for chemical reactions in real-world reactors, plays a crucial role in applications such as process monitoring, optimization, and advanced process control (Wu et al., 2019a; Hassanpour et al., 2020; Ellis and Chinde, 2020; Jalanko et al., 2021; Li et al., 2024a). Three idealized models are used to estimate the most important process variables of different chemical reactors: continuous stirred tank reactors (CSTRs), batch reactors (BRs), and plug flow reactors (PFRs). All three types of chemical reactors can be described using a set of ordinary differential equations (ODEs) or partial differential equations (PDEs), commonly referred to as first-principles models (Fogler, 2010). However, the parameters of these first-principles models vary depending on the specific reactions and reactors involved, and they can be difficult to determine accurately in real-world applications. Thus, there has been a transition to utilize data-driven methods, such as machine learning-based models, to capture the system dynamics of chemical reactions. These methods include least squares regression (Bhadriraju et al., 2019; Tang, 2023), feedforward neural networks (Liu et al., 2024), recurrent neural networks (RNNs) (Wu et al., 2019a, b), universal differential equation-based hybrid models (Bangi et al., 2022), physics-informed RNNs (PIRNNs) (Zheng et al., 2023), neural ordinary differential equations (Sorourifar et al., 2023), and input convex neural networks (Wang et al., 2025). However, a significant hurdle in employing deep learning within scientific domains is the scarcity of experimental data, which is normally costly and time-consuming to acquire via physical experimentation (e.g., it is resource-intensive to collect data and build neural network models for every new chemical reaction in the reactor). To mitigate this challenge, transfer learning-based approaches have been suggested (Xiao et al., 2023, 2024). However, for transfer learning-based methods, the process of identifying analogous reactions with adequate data, pre-training, and then transferring to the designated reaction must be repeated for each new reaction with a considerable number of samples, typically hundreds or thousands (Zhuang et al., 2020).
@@
This raises a fundamental question: can we develop a foundation model that rapidly adapts to any new reaction for chemical reactor modeling? To take a step toward this goal, we propose a model trained on a diverse dataset of integer-order chemical reactions in CSTRs, BRs, and PFRs. Firstly, we employ a meta-learning technique to build the model, specifically Reptile (Nichol and Schulman, 2018). After acquiring the model through Reptile, a conventional method to adapt the model to any new processes is via few-shot adaptation (the term “shot” refers to the number of data samples used to fine-tune the neural network). However, conventional data-driven adaptation in few-shot settings may result in poor generalization for nonlinear systems, due to the complexity of their state-space dynamics. To enhance the generalizability of the model and reduce the data required for adaptation, we draw inspiration from the success of physics-informed neural networks (Raissi et al., 2019). Specifically, we adopt a physics-informed adaptation approach by integrating domain knowledge, such as governing equations (e.g., mass and energy balances), into the adaptation process. Simulation data of hundreds of different reactions in three classical reactors is used to develop the model that will be further adapted to new reaction processes with limited data available. It is demonstrated that the adaptation process from the model to a specific reaction is efficient, requiring only a few shots for successful adaptation. The high-level system architecture is shown in Fig. 1 and the limitations of prior approaches and advantages of the proposed method are summarized in Fig. 2. In summary, the contributions of this work can be summarized in two main points:
  • 1.
    We introduce, for the first time in the literature, the concept of a foundation model specifically for chemical reactor modeling. This work serves as a significant first step toward developing a comprehensive foundation model in this field.
  • 2.
    We present meta-learning and physics-informed adaptation as key enabling techniques, demonstrating their effectiveness in facilitating rapid adaptation to unseen chemical reactors within a pool of ideal basic single-step chemical reactors (i.e., CSTRs, BRs, and PFRs).
Figure 1: High-level system architecture of “foundation model” for various chemical reactors.
Figure 2: Limitations of prior approaches and advantages of the proposed method.
@@
To avoid confusion and overstatement, we clarify that this work does not claim to fully develop a foundation model for chemical reactor modeling. Rather, it represents a step toward such models by proposing a framework in which meta-learning and physics-informed adaptation serve as key enablers. Accordingly, we use the term “foundation model” in quotation marks throughout the rest of this work to reflect this positioning. The rest of this article is structured as follows: Section 2 reviews the current works on physics-informed learning and meta-learning. Section 3 presents the algorithms for the Reptile-based “foundation model” with physics-informed adaptation. Section 4 outlines the formulation of CSTRs, BRs, and PFRs, covering reactor modeling and the design of physics-informed loss functions. Section 5 assesses the few-shot performance of the proposed method on unseen fixed-order CSTRs, BRs, and PFRs. Section 6 extends the method to incorporate random integer reaction orders. Finally, Section 7 addresses the limitations of the current design and proposes directions for future research.
@@
The effectiveness of transfer learning heavily relies on the similarity between the data distributions of the target and source tasks. In practice, finding a source reaction closely aligned with the target reaction is often challenging. A common approach to address this challenge involves conducting a few experiments under various conditions to gather data and develop a first-principles model through parameter optimization via regression. Once this model is established, it can be used to simulate a larger dataset, which then serves to pretrain a neural network for transfer learning, covering the entire operational range of the target reaction. However, transfer learning that depends on creating first-principles models to generate simulation datasets for each new reaction can be time-consuming, often taking hours or even days, and its effectiveness is highly contingent on the accuracy of the developed first-principles model. In contrast, our “foundation model” delivers high performance within seconds, eliminating the need for dependencies on first-principles models.
@@ 2 Meta-Learning and Physics-Informed Learning
Our goal and approach differ from these works. Due to the nature of our tasks, we use an RNN as the base model to develop a foundational model for generic chemical reactors (i.e., different parameterized ODE systems). Unlike conventional PINNs, which involve multiple applications of automatic differentiation, we fine-tune the meta-learned RNN using a physics-informed loss term computed with numerical methods, such as the forward Euler method, which requires fewer computational resources. Moreover, unlike foundation models in natural language processing (NLP) and computer vision (CV) - such as ChatGPT (Achiam et al., 2023) - which typically rely on self-supervised learning or reinforcement learning from human feedback (RLHF), our approach is grounded in meta-learning and physics-informed learning. These methods enable efficient adaptation to new chemical processes with limited data and are well-suited to the scientific modeling context. While our methodology differs from those used in large-scale language and vision models, it aligns with the core principles of foundation models: the ability to learn generalizable representations and adapt across tasks with minimal retraining. Although foundation models in NLP and CV are often evaluated across diverse task types, our model is designed to generalize across a wide range of reactor configurations and reaction systems within chemical process modeling. This eliminates the need to train separate models for each reactor type and supports broad adaptability within a relevant domain. Additionally, the integration of physics-informed adaptation ensures practical applicability in data-scarce settings, where zero-shot transfer may be unrealistic.
Furthermore, we believe that a successful foundation model should support both few-shot learning and continual learning, and existing literature indicates that meta-learning has significant potential for enabling continual learning (Riemer et al., 2018; Javed and White, 2019; Wang et al., 2024).
Remark 2.
While many foundation models - such as those described by Bommasani et al. (2021) - are large in scale, size alone does not define a foundation model. Its defining characteristics are generalization across diverse tasks, few-shot adaptability, and transferable representations. A central requirement is pretraining on a broad and diverse dataset, which we address by training on 1,500 simulated chemical systems encompassing CSTRs, BRs, and PFRs. Our model demonstrates scalability and extensibility by generalizing beyond its training distribution to unseen reaction kinetics, operating conditions, and reactor configurations. It is not limited to a single, narrowly defined task but adapts across multiple reactor systems without retraining. This aligns with emerging interpretations of foundation models in scoped, domain-relevant contexts. It should be noted that our model generalizes beyond specific training tasks to different reactors and conditions, consistent with the foundation model paradigm within a targeted domain. However, we do not claim generalization to a variety of process systems engineering tasks such as reactor design, fault detection, or control, as these tasks are fundamentally distinct due to task heterogeneity (e.g., classification vs. regression), differing mathematical structures (e.g., steady-state vs. dynamic models), and divergent modeling objectives. Importantly, while foundation models are expected to support diverse tasks, transferability is meaningful only when new tasks are aligned with the pretraining objective. Therefore, robust generalization within chemical reactor modeling represents a necessary and tractable first step.
Remark 3.
Rather than aiming to develop a universal foundation model for chemical engineering or process systems engineering, we position our model as a foundation model specifically for chemical reactor modeling, a coherent and practically significant subdomain. In this context, the downstream tasks correspond to different types of chemical reactors, each with distinct dynamics and kinetics. This scoping aligns with the framing adopted in recent literature, such as the work by Allen et al. (2024) on machine-learned interatomic potentials using meta-learning. In that study, the term “foundation model” is applied within the specific context of potential energy prediction, rather than across the entire domain of materials modeling. Similarly, we use the term in a domain-relevant manner to reflect our model’s ability to generalize across diverse reactor systems without task-specific retraining. Thus, our work contributes to the emerging application of foundation models in scientific machine learning, demonstrating how such models can be tailored to specialized domains while maintaining key properties of generalization and transferability.
@@ 3.1 Class of Systems
In this work, we consider chemical processes that can be described by a specific class of nonlinear ODEs, including those derived from PDEs using methods such as the method of lines:
@@ 3.2 Reptile for Meta-Learning
We begin by introducing the meta-learning algorithm used to build the “foundation model”. Specifically, we employ Reptile (Nichol and Schulman, 2018), a simplified version of MAML that prioritizes computational efficiency (Nichol et al., 2018). Unlike MAML, which relies on second-order derivatives, Reptile uses only first-order derivatives for meta-learning updates, making it more suitable for real-world applications. The primary goal of meta-training is to expose the model to a wide range of tasks, enabling it to learn initial parameters that facilitate rapid adaptation to new, unseen tasks. During this phase, Reptile optimizes the model’s parameters for efficient adaptation to different tasks with minimal updates. In essence, Reptile guides the model toward a solution $\theta$ that closely approximates the manifold of optimal solutions for each task $\tau$, by minimizing the Euclidean distance through repeated task sampling, model updates, and parameter adjustments. Mathematically, Reptile identifies network parameters $\theta$ that minimize the Euclidean distance $\mathcal{D}(\theta,\mathcal{W}_{\tau})$ for all tasks (i.e., locating a point near all solution manifolds), where $\mathcal{W}_{\tau}$ is the set of optimal parameters for task $\tau$:
@@ 3.3 Reptile with Physics-Informed Adaptation
Next, we develop a “foundation model” to learn the dynamics of various chemical reactions in three types of reactors. Our methodology comprises two distinct phases: meta-training and meta-testing. During meta-training (i.e., Algorithm 1), we immerse the model (i.e., specifically, an RNN) in a diverse array of chemical reactions across various reactor types. Specifically, the meta-training process involves two steps. First, we obtain the optimal parameters $\mathcal{W}_{\tau}$ for a specific task $\tau$ using the mean squared error (MSE) and the Adam optimizer (Kingma, 2014). Once $\mathcal{W}_{\tau}$ is determined, we update the network parameters $\theta$ using the Reptile update rule:
$\theta\leftarrow\theta+\alpha(\mathcal{W}_{\tau}-\theta)$ (3)
where $\alpha$ is the linear schedule step size hyperparameter. This process is then repeated recursively for all tasks.
@@
5:        Let $\mathcal{W}_{\tau}\leftarrow\theta$.
@@
9:           Update $\widetilde{\theta}\leftarrow Adam(\mathcal{L}_{\tau_{i}}(\widetilde{y_{i}},y_{i}),\mathcal{W}_{\tau})$.
@@
12:        Update $\theta\leftarrow\theta+\alpha(\mathcal{W}_{\tau}-\theta)$.
@@
Algorithm 1 Reptile-Based “foundation model” Meta-Training
In the subsequent meta-testing phase (i.e., Algorithm 2), we select previously unseen chemical reactions and undertake physics-informed few-shot adaptations to them, respectively, leveraging the Reptile-based “foundation model” obtained during meta-training. It should be pointed out that we use few-shot data $x$ (i.e., $x\in\mathcal{X}\subset D$, where $\mathcal{X}$ denotes the set of few-shot training data) to compute the data-driven loss term $\mathcal{L}_{d}$ and use collocation points $x_{c}$ (i.e., $x_{c}\in\mathcal{X}_{c}\subset D$, where $\mathcal{X}_{c}$ denotes the set of collocation data points, and $\mathcal{X}\cap\mathcal{X}_{c}=\emptyset$) to compute the physics-informed loss term $\mathcal{L}_{p}$ (a collocation point refers to a specific location within the domain of interest where the governing physics equations are enforced as part of the training process). As we do not utilize any labeled output data for $\mathcal{L}_{p}$, there is no need to perform physical experiments to gather additional data, which remains consistent with our few-shot setting. Moreover, it should be noted that we use estimated parameter values in $\mathcal{L}_{p}$ (i.e., fixed values) instead of their ground truth during adaptation (this differs from PIRNNs developed in our previous work (Zheng et al., 2023), which use ground truth parameter values), as the explicit parameter values of the first-principles models are not available in real-world scenarios. It is important to note that using approximate parameter values in physics-based regularization can introduce a discrepancy between the learned model and the true system dynamics. However, it is important to recognize that if we had perfect knowledge of physical laws and system parameters, a purely data-driven approach would be unnecessary. Physics-informed machine learning is particularly valuable for systems where data is limited, and domain knowledge is incomplete or imprecise. One of the key benefits of physics-informed machine learning is its ability to balance data-driven learning with physical constraints, even when the exact parameter values are uncertain. Numerous studies have demonstrated that even when physics-based constraints are not entirely accurate, incorporating them can improve model generalization and robustness compared to purely data-driven methods trained on limited data. In this work, the benefits of using physics-informed machine learning include:
  • 1.
    Rather than enforcing strict adherence to potentially incorrect kinetics, the regularization term serves as a soft constraints that incorporates domain knowledge. The model primarily learns from data, with the regularization acting as an additional informative bias rather than an absolute constraint.
  • 2.
    In the absence of exact kinetic parameters, the chosen sampling intervals are based on literature values and physical intuition. Taking the center points ensures that the regularization remains unbiased and provides a reasonable parameter estimation, preventing systematic skewing toward extreme values.
  • 3.
    Despite potential discrepancies, our results demonstrate that incorporating physics-based regularization improves the model’s generalizability and alignment with physical trends, validating its usefulness even with approximate parameter estimates.
The detailed formulation of physics-informed loss functions with respect to CSTRs, BRs, and PFRs will be presented in the following section, along with the parameterized first-principles models of CSTRs, BRs, and PFRs.
@@
Require: $M$: pre-trained “foundation model”
@@
We observed that employing the aforementioned approach to build a “foundation model” specifically for one type of reactor (e.g., CSTRs) can achieve a testing MSE on the order of $10^{-4}$ for normalized data easily with around 5 shots using a 2-layer RNN. However, based on our previous works (Wu et al., 2019a, b; Wang et al., 2025), we observed that a testing MSE on the order of $\mathbf{10^{-3}}$ for normalized data is indicative of satisfactory performance, rendering the model suitable for applications in process systems engineering such as neural network-based MPC.
@@
While developing a “foundation model” for each type of reactor might yield better performance for that specific reactor, a unified “foundation model” for all types of reactors could offer a significant advantage in generalizability. Specifically, it can adapt to the same chemical reaction across different reactor types, which a single-type “foundation model” cannot achieve. This adaptability is crucial when the real-world system involves the same reaction but in a different reactor, as a single-type model lacks information about other reactor types. This motivates us to develop a unified “foundation model” that can handle all three types of chemical reactors, rather than creating separate models for each type (our goal is not to develop an extremely precise model).
@@
4 Data Simulation for Chemical Reactors and Development of Physics-Informed Loss Term